can u help me with an equation plz the equation is:
x
- - 2=6
2

Found 2 solutions by fafa, Stargal510:
Answer by fafa(13)

(x/2)-2=6
(x/2)-2+2=6+2
(x/2)=8
(x/2).2=8.2
x=16

Answer by Stargal510(1)

um.... if ur solving for X, then u hav 2 get x by itself.
So first you Multiply both sides by 2.

x * 2 - 0 = 6
- * - - 2 = -
2 * 1 - 0 = 2
So now you have:
x - 2 = 3
Next, to find wat x equals you add both sides by 2
So it will turn out something like this :
x - 2 = 3
x + 2 = +2
----------
x = 5
So the answer would be X = 5
Hope this helped! =D